twelve times the sum of a number and 2 is equal to 36 what is the number
Answer by trx1150(14)

You can write each part of the problem algebraically, you just have to think of it as numbers and not as words.
"Twelve times the sum of a number and x" can be written as: 12(x+2)
When you see "a number" in a word problem, that will usually be your x value.
"...is equal to 36". Now we can write the whole equation as:
12(x+2)=36

To solve the equation, begin by distributing the 12 to both terms inside the parentheses.

12(x+2)=36
12x+24=36
Then, subtract 24 from each side.

12x=12
x=1
Hope this helped! :-)